# Diffusion-Driven Universal Model Inversion Attack for Face Recognition

### v1

@@ Abstract
Facial recognition technology poses significant privacy risks, as it relies on biometric data that is inherently sensitive and immutable if compromised. To mitigate these concerns, face recognition systems convert raw images into embeddings, traditionally considered privacy-preserving. However, model inversion attacks pose a significant privacy threat by reconstructing these private facial images, making them a crucial tool for evaluating the privacy risks of face recognition systems. Existing methods usually require training individual generators for each target model, a computationally expensive process. In this paper, we propose DiffUMI, a training-free diffusion-driven universal model inversion attack for face recognition systems. DiffUMI is the first approach to apply a diffusion model for unconditional image generation in model inversion. Unlike other methods, DiffUMI is universal, eliminating the need for training target-specific generators. It operates within a fixed framework and pretrained diffusion model while seamlessly adapting to diverse target identities and models. DiffUMI breaches privacy-preserving face recognition systems with state-of-the-art success, demonstrating that an unconditional diffusion model, coupled with optimized adversarial search, enables efficient and high-fidelity facial reconstruction. Additionally, we introduce a novel application of out-of-domain detection (OODD), marking the first use of model inversion to distinguish non-face inputs from face inputs based solely on embeddings.
I Introduction
I-A Background
@@
Face recognition technology presents significant privacy risks, as it involves processing biometric data that is both sensitive and immutable if compromised. Modern systems address these concerns by leveraging feature embedding techniques, which enhance scalability, generalization to unknown identities, and retrieval efficiency [1, 2, 3, 4]. These systems transform facial images into feature embeddings, facilitating recognition through distance-based metrics such as cosine similarity or Euclidean distance. This kind of approaches was traditionally considered privacy-preserving, as it encodes raw biometric data [5, 6], as shown in Fig. 1.
@@
Figure 1: Applications of our diffusion-driven universal model inversion (DiffUMI) for privacy attacks and out-of-domain detection (OODD). DiffUMI reconstructs images solely from embeddings, previously considered privacy-preserving, achieving 94.72% reconstruction success for face inputs, with reconstructions closely matching the target identity. Additionally, our OODD framework detects 98.9% of non-face inputs with a 3.9% error rate in misclassifying face inputs, demonstrating its robustness in distinguishing in-domain from out-of-domain data.
However, model inversion attacks present a significant privacy threat by reconstructing facial images solely from feature embeddings, as illustrated in Fig. 1. These reconstructions can facilitate further security breaches, including presentation attacks such as spoofing via photo, video replay, or 3D mask techniques [7]. As a result, model inversion attacks are essential for assessing the privacy vulnerabilities of face recognition systems [8].
@@
Generator Training-Free1 Input W/B2 Query3 Open-Set4 OODD5 Resolution6 Selfie7
@@
  • 1 “Training-Free” refers to methods where the generator is not target-specific, often pretrained on other tasks or not requiring a generator at all.
  • 2 White-box (W) model inversion attacks require access to the target model’s gradients, while black-box (B) attacks do not. “Both” indicates adaptability to both white-box and black-box settings.
  • 3 N/D: Information not disclosed. Note that for other attacks, W/B = Both yields the same query cost in both settings.
  • 4 Modern face recognition tasks inherently exhibit an “open-set” nature, allowing them to handle unknown identities that differ from the model’s training data. This requires attack methods capable of generalizing to an unbounded set of identities. Methods marked “No” are incompatible with open-set scenarios as they do not accept embeddings as inputs for model inversion.
  • 5 Open-set face recognition inherently accommodates out-of-domain inputs. We propose the first OODD framework for explicitly detecting such inputs based solely on embeddings through model inversion techniques.
  • 6 For example, “RGB160” refers to RGB images with a resolution of $160\times 160$ pixels. An absence of a number indicates undisclosed resolution.
  • 7 The reconstructed face varies in granularity, from isolated facial features to the full face or even the entire head (e.g., a selfie, highest fidelity).
  • 8 They define “OOD” as out-of-distribution, referring to a significant distributional shift between the face datasets used for training the target recognition model and the generator. This differs from our out-of-domain (OOD) definition, which describes instances such as a cat image in a face dataset.
  • 9 DiffUMI achieves comparable reconstruction accuracy in both white-box and black-box settings, though the higher computational cost in the black-box setting limits its practical applicability.
  • MAP2V is the most recent and relevant state-of-the-art benchmark aligned with our setting, both training-free and targeting open-set face recognition. Accordingly, we primarily compare against it under identical white-box conditions. Comparisons with attacks designed for closed-set face recognition are inapplicable due to differing assumptions: closed-set attacks rely on class labels, whereas we utilize target embeddings. Although DiffUMI also demonstrates superior performance over training-dependent attacks [22, 12] in Appendix -A, they are not our primary baseline due to their critical limitation that each target model requires a separately trained generator, incurring significant computational overhead.
@@
Finally, while diffusion models have gained prominence in modern generative AI research [27], generative adversarial networks (GANs) [28, 29, 30] remain the dominant paradigm for model inversion, as shown in Tab. I. This limited technological diversity restricts the application of state-of-the-art diffusion-driven generation techniques to model inversion. Existing diffusion-driven model inversion attacks typically rely on conditional diffusion (C-Diffusion) for facial image generation [19, 20], which are either inherently unsuitable for open-set face recognition or require significant computational resources, creating a trade-off between efficiency and image resolution (requiring 1.5 days [20] and 2 days [19] to train a generator restricted to $64\times 64$ pixel resolution). The closest training-free diffusion-driven approach to ours is PriDM [26]; however, it requires masked facial images as inputs, which contain significantly more information than embeddings alone. Therefore, we classify PriDM as a face inpainting method rather than a conventional model inversion technique.
I-B Motivation
@@
To address the aforementioned challenges, we propose DiffUMI, the first training-free, Diffusion-driven Universal Model Inversion attack. DiffUMI leverages a fixed, pretrained denoising diffusion model [31] for unconditional generation of full headshot-style selfies. Its universal nature arises from a consistent framework and generator that can adapt to any unknown identities and open-set face recognition models. DiffUMI eliminates the need for training target models or generators involving specific target identities, aligning with the open-set nature of face recognition and enhancing its suitability for open-set model inversion.
To achieve these objectives, we propose an algorithm that optimizes the diffusion model’s latent space via adversarial manipulation to refine reconstructions toward the target identity. We are the first to highlight the critical role of latent codes in generation fidelity and introduce an automated selection method for reliable latent codes from randomly sampled Gaussian distributions. Additionally, we identify the risk of overfitting to the target model when editing latent space and address this with a novel ranked adversary strategy. We further reveal that fine-grained adversarial attacks [32, 33] are most effective in optimizing latent space without introducing excessive distortions. However, we acknowledge that fine-grained black-box adversarial manipulation for high-resolution generation (e.g., $3\times 256\times 256$) is computationally expensive due to the high-dimensional latent space, posing a practical limitation for black-box diffusion-driven model inversion.
@@
We evaluate DiffUMI using two widely used face datasets [34, 35] and four face recognition models [36, 37, 5, 6], comparing its performance to three benchmarks [8, 22, 12]. The results highlight DiffUMI’s superior attack effectiveness, showing that an unconditional diffusion model with optimized adversarial search enables efficient, high-fidelity facial reconstruction. For example, when DiffUMI targets the Labeled Faces in the Wild (LFW) dataset [35] and the ArcFace model [37] in the white-box setting, it achieves Type I and Type II accuracy of 98.55% and 94.72%, surpassing the benchmarks by 9.57% and 15.5%, respectively. Furthermore, we successfully breach privacy-preserving face recognition models [5, 6], which were specifically designed to resist model inversion, achieving Type II accuracies ranging from 84.42% to 92.87% across two face datasets, surpassing the benchmarks by 4.01% to 9.82%. This raises critical concerns about the adequacy of existing privacy-preserving techniques in mitigating such threats.
@@
Moreover, we introduce a novel application of Out-Of-Domain Detection (OODD), marking the first use of model inversion to distinguish non-face inputs from face inputs based solely on feature embeddings. As shown in Fig. 1, the open-set nature of face recognition models enables non-face inputs, such as those from ImageNet [38], to be processed into embeddings, which we define as out-of-domain inputs. These embeddings share the same dimensionality and numerical range as genuine facial images, making differentiation inherently challenging. We propose a novel OODD framework utilizing model inversion techniques, where reconstructions of out-of-domain inputs typically fail either by not resembling the target identity or by lacking discernible human facial features. These failure cases serve as key indicators for identifying potential out-of-domain inputs. Our OODD framework effectively detects 98.9% of non-face inputs [38] after embedding by the ArcFace model [37], with only a 3.9% error rate for genuine face inputs [34, 35]. These findings demonstrate the potential of model inversion in robust OODD and offer a fresh perspective on the relationship between model inversion and feature space interpretability.
Our key contributions are summarized as follows:
  • We introduce DiffUMI, the first training-free, diffusion-driven universal model inversion attack, adaptable to any open-set face recognition model and unknown identities through a fixed framework and generator.
  • We present the first analysis of latent code influence on generation fidelity and propose an automated method for selecting highly reliable latent codes from randomly sampled Gaussian distributions.
  • We introduce a ranked adversary strategy to mitigate overfitting risks in latent space editing for generation. We demonstrate that fine-grained adversarial attack algorithms are particularly effective for manipulating the diffusion model’s latent space without introducing significant distortions or artifacts.
  • We empirically establish DiffUMI as a state-of-the-art attack, revealing critical vulnerabilities in privacy-preserving face recognition systems and raising concerns about their ability to counter such threats.
  • We introduce OODD, the first model inversion-based framework for distinguishing non-face inputs from face inputs based solely on feature embeddings.
@@
II Universal Model Inversion via Diffusion
This section introduces DiffUMI, outlining its objectives and framework in Sec. II-A. We then formalize the threat model and problem definition in Sec. II-B and Sec. II-C, respectively. Finally, we detail the algorithms for the three sequential steps in the framework, as presented in Sec. II-D through Sec. II-F. To make this paper self-contained, we outline the preliminaries of DDPM [39], face recognition, D’Agostino’s $K^{2}$ test [40, 41, 42], and MTCNN [43] in Appendix -B. A comprehensive notation list with corresponding definitions is provided in Tab. XIII of Appendix -C. The step-by-step algorithm of DiffUMI is detailed in Appendix -D.
II-A Overview
DiffUMI is introduced as a framework to assess the privacy risks of embedding-based face recognition models. A model is considered vulnerable if DiffUMI can reconstruct a facial image resembling the target identity using only its feature embedding. The primary goal of this privacy attack is to recover the target identity, rather than to precisely reconstruct the original image from which the embedding was derived.
@@ II-A Overview
  • Step (a) (Preparation) – Latent Code Generation: This phase independently generates a set of highly reliable latent codes, executed once, and applicable for attacking any target. The reliability of these latent codes is ensured by a selection strategy combining D’Agostino’s $K^{2}$ test [40, 41, 42] and MTCNN face detection [43]. The $K^{2}$ test ensures the latent codes follow a normal distribution, improving statistical reliability, while MTCNN ensures that the generated images contain explicit facial features. Together, these criteria enhance the latent codes’ robustness against subsequent adversarial manipulations, minimizing distortions and artifacts.
@@ II-B Threat Model
Attack Objective: Human assessment of privacy leakage, i.e., whether the reconstructed facial image resembles the target identity, is subjective and varies across individuals. Therefore, we evaluate the attack’s success using face recognition models, considering it successful if the reconstructed image is classified as the target identity.
Given that the attack process relies solely on the target embedding and model, there is a risk of overfitting due to adversarial artifacts rather than capturing true facial features. To mitigate this, we evaluate the attack’s success across multiple face recognition models, beyond just the target model. Formally, the attack objectives are as follows:
  • During attack: The reconstructed facial image’s embedding, when processed by the target model, achieves maximal similarity to the target embedding (see Sec. II-C).
  • During evaluation: The reconstructed facial image is classified as the target identity by multiple face recognition models, even when compared with other images of the same identity, excluding the one corresponding to the target embedding (see Sec. IV).
II-C Problem Definition and Objective Function
@@
II-D Step (a) - Preparation: Latent Code Generation
We propose a two-stage approach for generating reliable latent codes as candidate initializations for DiffUMI (Fig. 14 of Appendix -D). This strategy incorporates D’Agostino’s $K^{2}$ test [40, 41, 42] to ensure that the selected latent codes conform to a normal distribution, referred to as Gaussian normality, and utilizes MTCNN [43] for face detection to guarantee that the generated latent codes produce discernible facial features.
@@ II-D Step (a) - Preparation: Latent Code Generation
II-D1 Normality Test via D’Agostino’s K-Square Statistic
@@
Given a randomly generated latent code $x_{G}$, the $K^{2}$ test function $K(\cdot)$ [40, 41, 42] quantifies deviations from normality based on skewness and kurtosis, producing a probability value:
@@
where $\tau_{K}$ is chosen sufficiently large as generating a random Gaussian template remains computationally efficient. This selection process prioritizes latent codes closely adhering to Gaussian properties, enhancing robustness in subsequent steps.
II-D2 Face Detection via MTCNN

### v2

# DiffUMI: Training-Free Universal Model Inversion via Unconditional Diffusion for Face Recognition

@@ Abstract
Face recognition technology presents serious privacy risks due to its reliance on sensitive and immutable biometric data. To address these concerns, such systems typically convert raw facial images into embeddings, which are traditionally viewed as privacy-preserving. However, model inversion attacks challenge this assumption by reconstructing private facial images from embeddings, highlighting a critical vulnerability in face recognition systems. Most existing inversion methods require training a separate generator for each target model, making them computationally intensive. In this work, we introduce DiffUMI, a diffusion-based universal model inversion attack that requires no additional training. DiffUMI is the first approach to successfully leverage unconditional face generation without relying on model-specific generators. It surpasses state-of-the-art attacks by 15.5% and 9.82% in success rate on standard and privacy-preserving face recognition systems, respectively. Furthermore, we propose a novel use of out-of-domain detection (OODD), demonstrating for the first time that model inversion can differentiate between facial and non-facial embeddings using only the embedding space.
1 Introduction
@@
Face recognition technology presents significant privacy risks, as it involves processing biometric data that is both sensitive and immutable if compromised. Modern systems address these concerns by leveraging feature embedding techniques, which enhance scalability, generalization to unknown identities, and retrieval efficiency [1, 2, 3]. These systems convert facial images into feature embeddings, which are stored in a database and compared using distance-based metrics such as cosine similarity or Euclidean distance for recognition. This kind of approaches was traditionally considered privacy-preserving, as it encodes raw biometric data [4, 5, 6, 7], as shown in Fig. 1. However, model inversion attacks present a significant privacy threat by reconstructing facial images solely from feature embeddings. These reconstructions can facilitate further security breaches, including presentation attacks such as spoofing via photo, video replay, or 3D mask techniques [8]. As a result, model inversion attacks are essential for assessing the privacy vulnerabilities of face recognition systems [9].
@@
Figure 1: Our model inversion (DiffUMI) in privacy attacks [9] and out-of-domain detection (OODD) [10, 11, 12, 13]. DiffUMI reconstructs images solely from embeddings in the database, traditionally considered privacy-preserving, achieving a 94.72% reconstruction success rate for facial inputs, with outputs closely resembling the target identity. In parallel, our OODD framework detects 98.9% of non-face inputs while maintaining a low 3.9% false positive rate for face inputs, demonstrating strong capability in distinguishing in-domain from out-of-domain data.
@@
Generator Training-Free Input W/B Open-Set OODD Resolution Selfie
@@
  • denotes benchmark methods used for empirical comparison, representing the latest exemplars of each strategy. Comparisons with closed-set attacks and PriDM [31] are excluded due to fundamental differences in assumptions: they depend on class labels or images, while our approach operates on target embeddings.
@@
Finally, while diffusion models have gained prominence in modern generative AI research [32], generative adversarial networks (GANs) [33, 34, 35] remain the dominant paradigm for model inversion. This limited technological diversity restricts the application of state-of-the-art diffusion-driven generation techniques to model inversion. The most straightforward approach to diffusion-driven model inversion is to apply adversarial attacks (e.g., APGD [36]) directly in the latent space. However, this naive strategy produces significant artifacts due to the sensitivity of unconditional diffusion models, leading to overfitting and failed privacy recovery (see Appendix A). Consequently, existing approaches rely on conditional diffusion (C-Diffusion) for facial image generation [24, 25], which are either unsuitable for open-set face recognition or require substantial computational resources (taking 1.5 to 2 days to train generators restricted to $64\times 64$ resolution).
@@
To address the aforementioned challenges, we propose DiffUMI, the first training-free, Diffusion-driven Universal Model Inversion attack against embedding-based face recognition models. DiffUMI utilizes a fixed, pretrained denoising diffusion model [37] to unconditionally generate full headshot-style selfies, eliminating the need to train target-specific generators. Its universality stems from a consistent framework and generator capable of adapting to arbitrary identities and open-set face recognition models.
In terms of algorithm design, we hypothesize that if a reconstruction accurately matches the target identity without exhibiting adversarial artifacts, it successfully recovers private information. Thus, our objective is to maximize the embedding similarity between the reconstruction and the target while minimizing perceptual artifacts. Contrary to intuition, we found that the choice of attack backbone (e.g., APGD) is not the primary cause of artifacts. Instead, successful manipulation depends critically on three factors: reliable initialization of latent codes, fine-grained manipulation strategies, and mitigation of adversarial overfitting. To achieve these objectives, we first introduce an automated method for selecting reliable latent codes from randomly sampled Gaussian distributions. Second, we propose a ranked adversary strategy that performs fine-grained adversarial attacks [36, 38], guided by mathematically derived stopping criteria for optimizing the latent space in both white-box and black-box settings. To the best of our knowledge, our pipeline and algorithms are the first to effectively manipulate unconditional face generation.
Nevertheless, perfect artifact-free generation is inherently unattainable, and achieving it is not our goal. Instead, our objective is to recover the target identity from the embedding in a way that is visually recognizable to humans. The danger of adversarial artifacts lies in their potential to cause visually different images to be mapped to the same representation by the target model, leading to misleading matches. To eliminate the overfitting effect of these artifacts, which typically exhibit poor transferability, we evaluate privacy attacks using models different from the target. This evaluation protocol, widely adopted in prior work, assumes that if a reconstruction matches the target identity on models other than the one it was optimized against, it successfully recovers private information (a.k.a. transferability) [9, 17, 26, 27, 30].
@@
In practice, we assess DiffUMI on two widely used face datasets [39, 40] and four face recognition models [4, 5, 6, 7], comparing its performance against three state-of-the-art benchmarks [9, 27, 17]. In addition to automatic evaluation using face recognition models, we conduct user studies in which human participants are asked to judge whether the reconstruction matches the target identity or whether the target can be identified using the reconstruction alone. The results demonstrate DiffUMI’s strong attack effectiveness, showing that an unconditional diffusion model combined with optimized adversarial search enables efficient, high-fidelity facial reconstruction. For example, when targeting the Labeled Faces in the Wild (LFW) dataset [40] and the ArcFace model [5] in the white-box setting, DiffUMI achieves Type I and Type II accuracies of 98.55% and 94.72%, outperforming benchmarks by 9.57% and 15.5%, respectively. It also successfully breaches privacy-preserving face recognition models [6, 7], designed to resist model inversion, achieving Type II accuracies of 84.42% to 92.87% across two datasets, exceeding benchmarks by 4.01% to 9.82%. These findings raise serious concerns about the effectiveness of current privacy-preserving techniques.
@@
Moreover, we introduce a novel application of Out-Of-Domain Detection (OODD), marking the first use of model inversion to distinguish non-face from face inputs based solely on embeddings [10, 11, 12, 13]. As shown in Fig. 1, the open-set nature of face recognition models enables non-face inputs, such as those from ImageNet [41], to be processed into embeddings, which we define as out-of-domain inputs. These embeddings share the same dimensionality and numerical range as genuine facial images, making differentiation inherently challenging. Our OODD framework utilizes model inversion techniques, where reconstructions of out-of-domain inputs typically fail either by not resembling the target identity or by lacking discernible human facial features. These failure cases serve as key indicators for identifying potential out-of-domain inputs. Our OODD framework effectively detects 98.9% of non-face inputs [41] after embedding by the ArcFace model [5], with only a 3.9% error rate for genuine face inputs [39, 40].
The ability to distinguish real human face embeddings from out-of-domain inputs (e.g., animals or synthetic faces) can enhance face recognition systems by mitigating spoofing, deepfake enrollment, and data poisoning risks [10, 11, 12, 13]. It may support secure identity verification, clean large-scale face datasets, and ensure compliance with biometric standards. It may also aid in monitoring model robustness and controlling unintended use in applications like AR or photo filters. However, none of these works focus specifically on the face domain, embedding-level analysis, and open-set recognition, or leverage model inversion techniques.
Our key contributions are summarized as follows: (i) We introduce DiffUMI, the first training-free, diffusion-driven universal model inversion attack against embedding-based face recognition models. (ii) We propose the first algorithm that effectively manipulates unconditional face generation by automatically selecting highly reliable latent codes and introducing a novel ranked adversary strategy. (iii) We empirically establish DiffUMI as a state-of-the-art attack, revealing critical vulnerabilities in privacy-preserving face recognition systems and raising concerns about their ability to counter such threats. (iv) We introduce OODD, the first model inversion-based framework for distinguishing non-face inputs from face inputs based solely on feature embeddings.
@@
2 Universal Model Inversion via Diffusion
This section introduces DiffUMI, outlining its objectives and framework in Sec. 2.1. We then formalize the threat model and problem definition in Sec. 2.2 and Sec. 2.3, respectively. Finally, we detail the algorithms for the three sequential steps in the framework, as presented in Sec. 2.4 through Sec. 2.6. To make this paper self-contained, we outline the preliminaries of DDPM [42], face recognition, D’Agostino’s $K^{2}$ test [43, 44, 45], and MTCNN [46] in Appendix B. The step-by-step algorithm of DiffUMI and a comprehensive notation list are detailed in Appendix C.
2.1 Overview
DiffUMI is introduced as a framework to assess privacy risks of embedding-based face recognition models. A model is considered vulnerable if DiffUMI can reconstruct a facial image resembling the target identity using only its feature embedding. The primary goal of this privacy attack is to recover the target identity, rather than precisely reconstruct the original image from which the embedding was derived.
@@ 2.1 Overview
The core challenge addressed by DiffUMI is how to effectively manipulate unconditional diffusion generation. In the context of model inversion, this requires maximizing the embedding similarity between the target and the reconstruction while minimizing artifacts introduced by adversarial manipulation. While the similarity maximization is handled via a similarity-guided objective function (Sec. 2.3), the more difficult task of artifact minimization is tackled through our proposed three-step attack mechanism:
Step (a) (Preparation) – Latent Code Generation: This phase independently generates a set of highly reliable latent codes, executed once and applicable for attacking any target. The reliability is ensured by a selection strategy combining D’Agostino’s $K^{2}$ test [43, 44, 45] and MTCNN face detection [46]. The $K^{2}$ test ensures the latent codes follow a normal distribution, improving statistical reliability, while MTCNN ensures that the generated images contain explicit facial features. Together, these criteria enhance the latent codes’ robustness against subsequent adversarial manipulations, minimizing distortions and artifacts.
@@ 2.2 Threat Model
Attack Objective: The goal is to reconstruct a facial image that enables recognition of the target identity. To make this standard concrete and suitable for human evaluation, we define two criteria:
  • The reconstructed image appears to be the same person as the target.
  • The reconstructed image can help identify the target from a pool of identities.
However, human assessments require real-world user studies and are inherently subjective, varying across individuals. To provide a more objective measure, we also evaluate attack success using face recognition models. Since the attack relies solely on the target embedding and model, there is a risk of overfitting to adversarial artifacts rather than recovering genuine facial features. To mitigate this, we assess performance across multiple face recognition models beyond the target model. Formally, the attack objectives are:
  • During attack: Maximize the similarity between the reconstructed image’s embedding and the target embedding in the target model (see Sec. 2.3).
  • During evaluation: Ensure the reconstructed image is classified as the target identity by multiple models, even when compared against other images of the same person (excluding the one used for the target embedding) (see Sec. 4).
2.3 Problem Definition and Objective Function
@@
2.4 Step (a) - Prepare: Latent Code Generation
We propose a two-stage approach for generating reliable latent codes as candidate initializations for DiffUMI (Fig. 15 of Appendix C). This strategy incorporates D’Agostino’s $K^{2}$ test [43, 44, 45] to ensure that the selected latent codes conform to a normal distribution, referred to as Gaussian normality, and utilizes MTCNN [46] for face detection to guarantee that the generated latent codes produce discernible facial features.
@@ 2.4 Step (a) - Prepare: Latent Code Generation
2.4.1 Normality Test via D’Agostino’s K-Square Statistic
@@
Given a randomly generated latent code $x_{G}$, the $K^{2}$ test function $K(\cdot)$ quantifies deviations from normality based on skewness and kurtosis, producing a probability value:
@@
where $\tau_{K}$ is chosen sufficiently large as generating a random template remains computationally efficient. This selection process prioritizes latent codes closely adhering to Gaussian properties, enhancing robustness in subsequent steps.
2.4.2 Face Detection via MTCNN